OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

hello

pooja195 (pooja195):

\[y1-y2=m(x-x1)\] point slop formula now all you do is plug in these numbers

OpenStudy (anonymous):

can u help me some more were do i plug them in at

pooja195 (pooja195):

Y=mx+b slope=M

OpenStudy (jackofalltradez):

1 = y2 -7 = x2 -2/3 = m That should clarify things :)

OpenStudy (anonymous):

so the answer is ?

OpenStudy (jackofalltradez):

y + 7 = -2/3(x - 1)

OpenStudy (anonymous):

thx u so much